A fence post 5ft high casts a shadow 6ft long. At the same time the shadow of of a tower is 30 ft long. How high is the tower

Answer by alicealc(293)

tower's height = t
the ratio of height and shadow's length for the two objects should be the same, so:
5/6 = t/30
5 * 30 = 6t
150 = 6t
t = 150/6 = 25

so the tower is 25 ft high